How do you solve this problem?
Solve for L.
2L + 2W = P

Found 2 solutions by Abyssalanine, JBarnum:
Answer by Abyssalanine(3)

2L + 2W = P
2L = P - 2W
2L/2 = (P - 2W)/2
L=(P - 2W)/2

Answer by JBarnum(2146)